OpenStudy (anonymous):

What is the solution for this system of linear equations? y=-3x-5 y=-8x-25 so -8x-25=-3x+5

OpenStudy (anonymous):

@jim_thompson5910 can you work through this with me? do I subtract -3 or add it?

OpenStudy (anonymous):

-8x +3x = 5 + 25

OpenStudy (anonymous):

now it's arithmetic

jimthompson5910 (jim_thompson5910):

y=-3x-5 -8x-25=-3x-5 ... plug in y = -8x-25 If you want to move the -3x over, you add 3x to both sides to get -8x+3x-25 = -3x+3x - 5 -5x - 25 = -5

OpenStudy (anonymous):

:) it's great ur asking for someone to work it with u rather than just wanting the answer! :)

jimthompson5910 (jim_thompson5910):

so what's next after you get -5x - 25 = -5

OpenStudy (anonymous):

subtract -5 from -25?

jimthompson5910 (jim_thompson5910):

how about adding 25 to both sides -5x - 25 = -5 -5x - 25+25 = -5+25 -5x +0 = 20 -5x = 20 We do this because we want to move the -25 over

OpenStudy (anonymous):

okay, so x=-4?

jimthompson5910 (jim_thompson5910):

good

jimthompson5910 (jim_thompson5910):

and you know what to do to get y

OpenStudy (anonymous):

and y=7? :)

jimthompson5910 (jim_thompson5910):

yep x = -4 y = 7

OpenStudy (anonymous):

@HawkCrimson :)

OpenStudy (anonymous):

thank you @jim_thompson5910 !! :)

jimthompson5910 (jim_thompson5910):

you're welcome

OpenStudy (anonymous):

:)